Calculate the energy present in one mole of quanta of Indio light with wavelength 430 nm
1
Expert's answer
2020-10-21T13:19:44-0400

"E=\frac{hc}{λ}=\frac{6.626\times10^{-34}\times 2.998\times 10^8}{430\times10^{-9}}=4.62\times10^{-19} J"

in one mole:

"E(1 mole)=4.62\times10^{-19}J \times 6.022\times10^{23}= 2.78\times10^5 J"